# The Rise of Small Language Models in Healthcare: A Comprehensive Survey

### v1

@@ 1. Background
The global small language model (SLM) market is projected to reach USD 29.64 billion by 2032, with the healthcare industry anticipated to experience the fastest compound annual growth rate (CAGR) of 18.31% from 2024 to 2032222https://www.globenewswire.com/. This growth is driven by the growing adoption of SLM in medical diagnosis, patient care, and administrative processes. The use of SLM in analyzing extensive medical literature and patient data provides clinicians with evidence-based interpretations to support clinical decision making [^3] https://www.healthcare.digital/.
@@ 1.1. Motivation
The deployment of language models in clinical settings demands an uncertain trade-off among high-stakes accuracy, regulatory compliance, and computational efficiency (Singh et al., 2023). While LLMs effectively capture clinical concepts (Fu et al., 2023), their high inference latency (>500ms per query) and memory demands (>16GB VRAM) make them unsuitable for real-time analysis of medical records with protected health information (PHI). Healthcare data is governed by stringent regulations such as HIPAA and GDPR and requires on-premise processing to mitigate privacy risks (Jonnagaddala and Wong, 2025). Regulatory agencies (e.g., FDA [^4] https://www.fda.gov/) encourages transparent AI-driven decision-making to promote interpretation. Knowledge-driven models generate interpretable results based on trusted external medical knowledge bases, supporting regulatory compliance such as EU Medical Device Regulation (MDR) (Gilbert, 2023). Training LLMs cause an immense carbon footprint of more than 500 metric tons $\text{CO}_{2}$ which is not sustainable for environmental goals. In contrast, SLM is optimized for watts per inference, a sustainability metric (Potkins et al., 2024). They reduce energy consumption by 10 to 100 times through optimization and compression techniques, addressing the Jevons paradox in green AI (making models more energy-efficient can sometimes lead to an overall increase in resource consumption—not a decrease, as one might expect) (Luccioni et al., 2025). Critical care scenarios require subsecond decision making driven by clinical NLP (cNLP) tasks where LLMs with high inference latencies are not suitable for time-sensitive applications (Bai et al., 2024). As such, we opt to expose the potential of SLM to address specialized healthcare challenges under limited resources and time constraints.
@@ 1.2. Language model fundamentals
Transformers overview: SLMs are transformer-based models that capture contextual relationships and dependencies in language. A transformer is a deep learning model that processes sequences in parallel by converting tokenized words into numerical embeddings. Transformers use positional encoding to maintain word order and allow parallel processing. The transformer encoder converts input sequences into context-rich representations, while the decoder generates output sequences. Transformers are categorized into encoder-only, decoder-only, and encoder-decoder architectures. The encoder captures relationships among words using self-attention mechanisms in the attention layer, assigning weights based on contextual relevance (Belhaouari and Kraidia, 2025; Liu and Xu, 2023). Encoder-only models (BERT-like) are generally good in discriminative tasks. Decoder-only models (GPT-like) generate text by predicting the next token in a sequence, making them ideal for generative tasks. Encoder-decoder models (e.g., T5, BART (Lewis et al., 2019)) combine input understanding and output generation. Given our focus on generative tasks in clinical settings, we emphasize decoder-only SLMs for their efficiency and adaptability.
@@ 1.3. Scope
Early studies define SLM as the models with 1B to 20B parameters, while LLMs have hundreds of billion parameters (Subramanian et al., 2025). Industry research defines SLMs based on deployment limits, such as models running on CPU-only edge devices (Sun et al., 2020). Later research suggests a relative definition, where SLMs are 10 to 100 times smaller than the largest models of their time (Jovanovic and Campbell, 2024). The definition of “small" in SLMs is relative in time and evolves with hardware advancements. It generally refers to reduced computational complexity (e.g., parameter count) and lower storage requirements (e.g., model size).
@@
Search strategy: The existing literature used mixed terminology for referring to the SLM: local language models, remote language models, compressed models, optimized models, on-device language models, small LLMs and lightweight language models. We used these keywords to identify SLMs designed for NLP-driven clinical decision making from well-known sources: PubMed, ACM, IEEE, Google Scholar, Scopus. Our study covers three dimensions: (i) NLP tasks, (ii) stakeholders, (iii) care continuum. The NLP tasks covers question answering (QA), named entity recognition (NER), relation extraction (RE), information extraction (IE), natural language inference/ understanding (NLI/ NLU), text classification (CLS), summarization (SUM). We focus on three important stakeholders: patient, healthcare professionals, research/material. The care continuum covers triaging, diagnosis, treatment, recovery, risk prediction, and virtual health assistants.
Inclusion criteria. Unlike fundamental bidirectional BERT-like transformers, our interest lies in the decoder-only transformer architecture models (traced back to GPT) in the healthcare domain with open weights. We include SLMs trained on English language data for clinical NLP tasks, developed before April 2025. The scope of our study is defined in Fig. 1(a). Exclusion criteria. We exclude models that are non-English, designed for inferring non-textual audiovisual data, multimodal SLMs, agentic AI systems, and models exceeding 7 billion parameters. Additionally, models lacking peer-reviewed publications for SLM models are not considered in this study.
@@ 1.4. Contributions
    Outline the role of SLMs in healthcare applications (§5.2), examine associated challenges, and highlight key research directions for clinical integration (§6).
@@ 2. Architectural foundations: Building SLMs from scratch
The lightweight architectures are designed for efficient performance with fewer parameters and reduced computational overhead while maintaining high performance in medical applications. While model efficiency is important, the success of SLMs in healthcare also depends on the quality and accessibility of medical data. This motivates a closer look at the data landscape in clinical settings. In this section, we identify SLMs in healthcare, their timeline, and configurations (§2.1) to focus on architectural foundations in healthcare through healthcare data (§2.2), pre-training components (§2.3), and attention mechanisms (§2.4).
@@ 2.1. SLM in healthcare: Timeline and configuration
Figure 3. Timeline of SLM development in healthcare. Transforming NLP-driven clinical decision-making, NLP tasks for medical conditions, and doctor-patient conversation or mental health analysis.
We present the timeline of model development in Fig. 3. SLMs are configured with different aspects to optimize their performance. Notably, few studies provide detailed configurations of SLMs for healthcare applications. Table 2 presents the available configurations we identified, offering an overview of current development efforts. The model size, measured in parameters, typically ranges from 124 million to 7 billion weights. The model depth, defined by the number of transformer layers, typically spans from 22 to 32 layers. The hidden dimension, or the size of the model’s embedding vectors, usually ranges from 1024 to 4096 units. To optimize attention processing, SLMs employ 16 to 64 parallel attention heads, enabling the model to focus on various aspects of the data simultaneously. Additionally, the context window defines the maximum sequence length, allowing the model to process up to 4096 tokens, which is crucial for handling long clinical texts such as patient records and discharge summaries. These design choices ensure that SLMs achieve both high accuracy and efficiency in healthcare applications.
@@ 2.2. Healthcare data
The datasets used for SLM construction are shown in Table 4. Yan et al. (2024) highlights the challenges with data privacy and the lack of the open domain datasets. A potential solution is the synthetic health data generation (SHDG), leveraging learning from statistical distributions of real-world datasets. Recent reviews and surveys on SHDG (Gonzales et al., 2023; Pezoulas et al., 2024) provide a quick reference guide for selecting appropriate synthetic data generation strategies. Recent studies used LLMs for SHDG and validated that LLM-generated synthetic data can improve system performance (Kang et al., 2024). A comprehensive study on metrics for SHDG evaluation (Osorio-Marulanda et al., 2024) noted the absence of a composite metric for quality assessment. SHDG supports data enhancement888augmenting, interpolating, or fusing with synthetic data and de-identification999masking personally identifiable information from patient records in healthcare but also raises concerns about vulnerabilities such as data poisoning attacks101010malicious data is inserted into the training set, intentionally corrupting healthcare models to produce inaccurate or biased outputs that can compromise model integrity. Data poisoning attack (Zhao et al., 2025) occurs when even 0.001% of training tokens are replaced by medical misinformation (Alber et al., 2025). Lobo et al. (2022) emphasizes the need for robust evaluation against data poisoning, which can be mitigated through secure data practices, thorough cleaning, and multistage model training (He et al., 2025).
@@
Table 3. Comparison of healthcare SLMs: Year, training/evaluation datasets, and data scale.
@@
Data cleaning mechanisms are de-identification, de-sensitization, and de-duplication, that enhance model performance by reducing redundancy and increasing diversity. De-duplication improves data quality and prevents data leakage between training and test sets, as demonstrated by models like MediTron (Sallinen et al., 2025) and Apollo (Wang et al., 2024a). Importance resampling efficiently selects relevant data in a reduced feature space, outperforming existing methods and improving downstream model accuracy (Xie et al., 2023a). Notably, Shi et al. (2024) employ perplexity and k-center greedy methods to increase data quality and data diversity, respectively, while developing SLM. After surveying data diversity metrics, Yang et al. (2025) proposed an effective data selection strategy; however, its application in the healthcare domain remains unexplored, underscoring the need to benchmark selection methods for medical datasets to enhance diversity, minimize redundancy, and prevent overfitting. We enlist the training and evaluation datasets used in the existing SLMs for healthcare in Table 3 and mention open source datasets for each NLP task in Table 4.
@@ 2.4. Attention mechanisms
Core attention architectures: Core attention mechanisms in language models can be optimized for healthcare by investigating how clinical concepts and contextual information interact during processing. The attention mechanism has evolved from self-attention to cross-attention, and further to multi-head, entity-aware, and knowledge-guided variants — each contributing to advancements in healthcare applications. For example, Self-attention is useful in attending certain medical aspect— symptoms, diagnoses, or medications, capturing dependencies like the relationship between “chest pain" and “ECG findings" (Vaswani et al., 2017). Cross-attention enables one source of information (e.g., lab results) to attend to another (e.g., physician notes), supporting rich understanding by integrating information from both structured and unstructured information (Lai et al., 2024). Global attention highlights key clinical anchors (e.g., “principal diagnosis” or “chief complaint”) that influence interpretation of the entire record (Liu, 2025). Multi-head attention (MHA) models multiple types of clinical relationships in parallel-temporal patterns, medication-disease interactions, or symptom clusters (Meng et al., 2025). The knowledge-guided attention integrates biomedical ontologies or knowledge graphs (e.g., UMLS) to bias attention toward medically validated relationships, enhancing clinical accuracy and explainability (Li et al., 2023b). However, a key limitation of core attention mechanisms—particularly self-attention—is their memory inefficiency, as the computational and memory cost scales quadratically with the input length, making it challenging to process long clinical documents or patient histories in EHR.
@@
Efficient attention: To address computational bottlenecks, SLM in healthcare leverage memory-efficient attention adaptations. Sparse attention restricts these interactions to the most relevant information— focusing only on recent or abnormal findings—making it ideal for processing lengthy medical notes efficiently (Beltagy et al., 2020; Child et al., 2019). Multi-query attention (MQA) allows multiple attention heads to share the same keys and values (KV) pairs (Xu et al., 2023). For example, in a clinical note, different attention heads can be used to extract aspects like demographic information, medication history, and clinical conditions, while sharing the same KV pairs. Grouped-query attention (GQA) groups query heads into clusters, with each cluster sharing a KV pair (Ainslie et al., 2023). For example, when extracting SDoH, one cluster of heads might focus on housing, another on employment, and another on education. Multi-Head Latent Attention uses low-rank compression of KV pairs to reduce the need for a KV cache (Lu et al., 2024a). For example, a model extracting QoL measures from patient feedback can use this compression to handle large datasets without storing full-dimensional attention matrices for each concept. QCQA introduces a quality- and capacity-aware grouped query attention mechanism that optimally balances key-value (KV)-cache size and text generation accuracy, outperforming GQA by up to 20% in accuracy and reducing memory requirements by 40% without fine-tuning (Joshi et al., 2024). Flash Attention (FA) divides the attention matrix into smaller blocks and recomputes them during backpropagation, saving up to 4 times more memory without losing precision (Dao et al., 2022). For example, in a clinical question-answering system, instead of storing the entire attention matrix for a patient’s full medical report, FA stores only key blocks and recomputes others as needed, reducing memory usage and allowing the system to handle larger datasets like multi-visit histories.
@@
Low-Rank Factorization approximates the full attention matrix by decomposing it into smaller, lower-dimensional matrices, retaining only the essential relationships. In a clinical concept extraction model, this approach increases efficiency by replacing large KV matrices with more compact representations. A series of advancements in low-rank adaptation techniques have emerged to optimize LLM fine-tuning. Memory-Efficient Attention optimizes processing by calculating attention scores without storing the entire matrix. In a real-time clinical decision support system, this approach allows for dynamic attention score calculation using medical records, enabling efficient analysis of patient symptoms and medical history without storing full context. Sliding Window Attention captures local dependencies by focusing on a fixed-sized window of neighboring tokens. The clinical note SUM allows the model to focus on specific sections such as the history of the present illness. Random Projection Attention uses dimensionality reduction to approximate softmax attention, improving scalability. Building on these foundations, recent advancements have given rise to more sophisticated architectures. Notably, Mamba (Gu and Dao, 2023) introduces a selective state space model characterized by input-dependent transitions. In predictive models for patient outcomes, it reduces the complexity of processing high-dimensional data, such as medical tests and demographics, making the model more efficient and computationally cost-effective for large datasets and real-time predictions. We analyze the time complexity of each attention mechanism to shed some light on how healthcare research can integrate attention-driven component based on available resources in Table 5.
@@
Table 5. Time complexities of various attention mechanisms in clinical NLP, from least to most efficient. Here, $N$ is the sequence length, $d$ is the embedding dimension, $M$ is the cross-attended length, $w$ is a local window size, $g$ is the number of global tokens, and $r$ is a low-rank dimension. “KG overhead” and $\epsilon$ indicate extra constant factors due to knowledge bases or entity-related modules. These values are based on official reports or estimated according to functionality.

### v2

@@ 1. Background
The global small language model (SLM) market is projected to reach USD 29.64 billion by 2032, with the healthcare industry anticipated to experience the fastest compound annual growth rate (CAGR) of 18.31% from 2024 to 2032222https://www.globenewswire.com/. This growth is driven by the growing adoption of SLM in medical diagnosis, patient care, and administrative processes. The use of SLM in analyzing extensive medical literature and patient data provides clinicians with evidence-based interpretations to support clinical decision support system (CDSS) [^3] https://www.healthcare.digital/.
@@ 1.1. Motivation
The deployment of language models in hospital settings demands an uncertain trade-off among high-stakes accuracy, regulatory compliance, and computational efficiency (Singh et al., 2023). While LLMs effectively capture clinical concepts (Fu et al., 2023), their high inference latency (>500ms per query) and memory demands (>16GB VRAM) make them unsuitable for real-time analysis of medical records with protected health information (PHI). Healthcare data is governed by stringent regulations such as HIPAA and GDPR and requires on-premise processing to mitigate privacy risks (Jonnagaddala and Wong, 2025). Regulatory agencies (e.g., FDA [^4] https://www.fda.gov/) encourages transparent AI-driven decision-making to promote interpretation. Knowledge-driven models generate interpretable results based on trusted external medical knowledge bases, supporting regulatory compliance such as EU Medical Device Regulation (MDR) (Gilbert, 2023). Training LLMs cause an immense carbon footprint of more than 500 metric tons $\text{CO}_{2}$ which is not sustainable for environmental goals. In contrast, SLM is optimized for watts per inference, a sustainability metric (Potkins et al., 2024). They reduce energy consumption by 10 to 100 times through optimization and compression techniques, addressing the Jevons paradox in green AI (making models more energy-efficient can sometimes lead to an overall increase in resource consumption—not a decrease, as one might expect) (Luccioni et al., 2025). Critical care scenarios require subsecond decision making driven by clinical NLP (cNLP) tasks where LLMs with high inference latencies are not suitable for time-sensitive applications (Bai et al., 2024). As such, we opt to expose the potential of SLM to address specialized healthcare challenges under limited resources and time constraints.
@@ 1.2. Language model fundamentals
Transformers overview: SLMs are transformer-based models that capture contextual relationships and dependencies in language. A transformer is a deep learning model that processes sequences in parallel by converting tokenized words into numerical embeddings. Transformers use positional encoding to maintain word order and allow parallel processing. The transformer encoder converts input sequences into context-rich representations, while the decoder generates output sequences. Transformers are categorized into encoder-only, decoder-only, and encoder-decoder architectures. The encoder captures relationships among words using self-attention mechanisms in the attention layer, assigning weights based on contextual relevance (Belhaouari and Kraidia, 2025; Liu and Xu, 2023). Encoder-only models (BERT-like) are generally good in discriminative tasks. Decoder-only models (GPT-like) generate text by predicting the next token in a sequence, making them ideal for generative tasks. Encoder-decoder models (e.g., BART (Lewis et al., 2019)) combine input understanding and output generation. Given our focus on generative tasks in hospital settings, we emphasize decoder-only SLMs for their efficiency and adaptability.
@@ 1.3. Scope
Early studies define SLM as the models with 1B to 20B parameters, while LLMs have hundreds of billion parameters (Subramanian et al., 2025). Industry research defines SLMs based on deployment limits, such as models running on CPU-only edge devices (Sun et al., 2020). Later research suggests a relative definition, where SLMs are 10 to 100 times smaller than the largest models of their time (Jovanovic and Campbell, 2024). The definition of “small” in SLMs is relative in time and evolves with hardware advancements. It generally refers to reduced computational complexity (e.g., parameter count) and lower storage requirements (e.g., model size).
@@
Search strategy: The existing literature used mixed terminology for referring to the SLM: local language models, remote language models, compressed models, optimized models, on-device language models, small LLMs and lightweight language models. We used these keywords to identify SLMs designed for NLP-driven CDSS from well-known sources: PubMed, ACM, IEEE, Google Scholar, Scopus. Our study covers three dimensions: (i) NLP tasks, (ii) stakeholders, (iii) care continuum. The NLP tasks covers question answering (QA), named entity recognition (NER), relation extraction (RE), information extraction (IE), natural language inference/ understanding (NLI/ NLU), text classification (CLS), summarization (SUM). We focus on three important stakeholders: patient, healthcare professionals, research/material. The care continuum covers triaging, diagnosis, treatment, recovery, risk prediction, and virtual health assistants.
Inclusion criteria. Unlike fundamental bidirectional BERT-like transformers, our interest lies in the decoder-only transformer architecture models (traced back to GPT) in the healthcare domain with open weights. We include SLMs trained on English language data for cNLP tasks, developed before April 2025. The scope of our study is defined in Fig. 1(a). Exclusion criteria. We exclude models that are non-English, designed for inferring non-textual audiovisual data, multimodal SLMs, agentic AI systems, and models exceeding 7 billion parameters. Additionally, models lacking peer-reviewed publications for SLM models are not considered in this study.
@@ 1.4. Contributions
    Outline the role of SLMs in healthcare applications (§5.2), examine associated challenges, and highlight key research directions for healthcare integration (§6).
@@ 2. Architectural foundations: Building SLMs from scratch
The lightweight architectures are designed for efficient performance with fewer parameters and reduced computational overhead while maintaining high performance in medical applications. While model efficiency is important, the success of SLMs in healthcare also depends on the quality and accessibility of medical data. This motivates a closer look at the data landscape in hospital settings. In this section, we identify SLMs in healthcare, their timeline, and configurations (§2.1) to focus on architectural foundations in healthcare through healthcare data (§2.2), pre-training components (§2.3), and attention mechanisms (§2.4).
@@ 2.1. SLM in healthcare: Timeline and configuration
Figure 3. Timeline of SLM development in healthcare. Transforming NLP-driven clinical decision-making, NLP tasks for clinical conditions, and doctor-patient conversation or mental health analysis.
We identified 20 SLMs in the literature based on our search criteria and selection principles. We present the timeline of model development in Fig. 3. SLMs are configured with different aspects to optimize their performance. Notably, very few (4 out of 20) studies provide detailed configurations of SLMs for healthcare applications. Table 2 presents the available configurations we identified, offering an overview of current development efforts. The model size, measured in parameters, typically ranges from 124 million to 7 billion weights. The model depth, defined by the number of transformer layers, typically spans from 22 to 32 layers. The hidden dimension, or the size of the model’s embedding vectors, usually ranges from 1024 to 4096 units. To optimize attention processing, SLMs employ 16 to 64 parallel attention heads, enabling the model to focus on various aspects of the data simultaneously. Additionally, the context window defines the maximum sequence length, allowing the model to process up to 4096 tokens, which is crucial for handling long clinical texts such as patient records and discharge summaries. These design choices ensure that SLMs achieve both high accuracy and efficiency in healthcare applications.
@@ 2.2. Healthcare data
The datasets used for SLM construction are shown in Table 4. Yan et al. (2024) highlights the challenges with data privacy and lack of the open domain datasets. A potential solution is the synthetic health data generation (SHDG), leveraging learning from statistical distributions of real-world datasets. Recent reviews and surveys on SHDG (Gonzales et al., 2023; Pezoulas et al., 2024) provide a quick reference guide for selecting appropriate synthetic data generation strategies. Recent studies used LLMs for SHDG and validated that LLM-generated synthetic data can improve system performance (Kang et al., 2024). A comprehensive study on metrics for SHDG evaluation (Osorio-Marulanda et al., 2024) noted the absence of a composite metric for quality assessment. SHDG supports data enhancement888augmenting, interpolating, or fusing with synthetic data and de-identification999masking personally identifiable information from patient records in healthcare but also raises concerns about vulnerabilities such as data poisoning attacks101010malicious data is inserted into the training set, intentionally corrupting healthcare models to produce inaccurate or biased outputs that can compromise model integrity. Data poisoning attack (Zhao et al., 2025) occurs when even 0.001% of training tokens are replaced by medical misinformation (Alber et al., 2025). Lobo et al. (2022) emphasized the need for robust evaluation against data poisoning, which can be mitigated through secure data practices, thorough cleaning, and multistage model training (He et al., 2025).
@@
Table 3. Comparison of healthcare SLMs: Year and training/evaluation datasets.
@@
Data cleaning mechanisms are de-identification, de-sensitization, and de-duplication, that enhance model performance by reducing redundancy and increasing diversity. De-duplication improves data quality and prevents data leakage between training and test sets, as demonstrated by models like MediTron (Sallinen et al., 2025) and Apollo (Wang et al., 2024a). Importance resampling efficiently selects relevant data in a reduced feature space, outperforming existing methods and improving downstream model accuracy (Xie et al., 2023a). Notably, Shi et al. (2024) employed perplexity and k-center greedy methods to increase data quality and data diversity, respectively, while developing SLM. After surveying data diversity metrics, Yang et al. (2025) proposed an effective data selection strategy; however, its application in the healthcare domain remains unexplored, underscoring the need to benchmark selection methods for medical datasets to enhance diversity, minimize redundancy, and prevent overfitting. We enlist the training and evaluation datasets used in the existing SLMs for healthcare in Table 3 and mention open source datasets for each of the NLP task in healthcare (see Table 4).
@@ 2.4. Attention mechanisms
Core attention architectures: Core attention mechanisms in language models can be optimized for healthcare by investigating how clinical concepts and contextual information interact during processing. The attention mechanism has evolved from self-attention to cross-attention, and further to multi-head, entity-aware, and knowledge-guided variants — each contributing to advancements in healthcare applications. For example, Self-attention is useful in attending certain medical aspect— symptoms, diagnoses, or medications, capturing dependencies like the relationship between “chest pain” and “ECG findings” (Vaswani et al., 2017). Cross-attention enables one source of information (e.g., lab results) to attend to another (e.g., physician notes), supporting rich understanding by integrating information from both structured and unstructured information (Lai et al., 2024). Global attention highlights key clinical anchors (e.g., “principal diagnosis” or “chief complaint”) that influence interpretation of the entire record (Liu, 2025). Multi-head attention (MHA) models multiple types of clinical relationships in parallel-temporal patterns, medication-disease interactions, or symptom clusters (Meng et al., 2025). The knowledge-guided attention integrates biomedical ontologies or knowledge graphs (e.g., UMLS) to bias attention toward medically validated relationships, enhancing clinical accuracy and explainability (Li et al., 2023b). However, a key limitation of core attention mechanisms—particularly self-attention—is their memory inefficiency, as the computational and memory cost scales quadratically with the input length, making it challenging to process long healthcare-related documents and patient histories in EHR.
@@
Efficient attention: To address computational bottlenecks, SLM in healthcare leverage memory-efficient attention adaptations. Sparse attention restricts these interactions to the most relevant information— focusing only on recent or abnormal findings—making it ideal for processing lengthy medical notes efficiently (Beltagy et al., 2020; Child et al., 2019). Multi-query attention (MQA) allows multiple attention heads to share the same keys and values (KV) pairs (Xu et al., 2023). For example, in a clinical note, different attention heads can be used to extract aspects like demographic information, medication history, and clinical conditions, while sharing the same KV pairs. Grouped-query attention (GQA) groups query heads into clusters, with each cluster sharing a KV pair (Ainslie et al., 2023). For example, when extracting social determinants of health (SDoH), one cluster of heads might focus on housing, another on employment, and another on education. Multi-Head Latent Attention uses low-rank compression of KV pairs to reduce the need for a KV cache (Lu et al., 2024a). For example, a model extracting quality of life (QoL) measures from patient feedback can use this compression to handle large datasets without storing full-dimensional attention matrices for each concept. QCQA introduces a quality- and capacity-aware grouped query attention mechanism that optimally balances key-value (KV)-cache size and text generation accuracy, outperforming GQA by up to 20% in accuracy and reducing memory requirements by 40% without fine-tuning (Joshi et al., 2024). Flash Attention (FA) divides the attention matrix into smaller blocks and recomputes them during backpropagation, saving up to 4 times more memory without losing precision (Dao et al., 2022). For example, in a clinical question-answering system, instead of storing the entire attention matrix for a patient’s full medical report, FA stores only key blocks and recomputes others as needed, reducing memory usage and allowing the system to handle larger datasets like multi-visit histories.
@@
Low-Rank Factorization approximates the full attention matrix by decomposing it into smaller, lower-dimensional matrices, retaining only the essential relationships. In a clinical concept extraction model, this approach increases efficiency by replacing large KV matrices with more compact representations. A series of advancements in low-rank adaptation techniques have emerged to optimize LLM fine-tuning. Memory-Efficient Attention optimizes processing by calculating attention scores without storing the entire matrix. In a real-time CDSS, this approach allows for dynamic attention score calculation using medical records, enabling efficient analysis of patient symptoms and medical history without storing full context. Sliding Window Attention captures local dependencies by focusing on a fixed-sized window of neighboring tokens. The clinical note SUM allows the model to focus on specific sections such as the history of the present illness. Random Projection Attention uses dimensionality reduction to approximate softmax attention, improving scalability. Building on these foundations, recent advancements have given rise to more sophisticated architectures. Notably, Mamba (Gu and Dao, 2023) introduces a selective state space model characterized by input-dependent transitions. In predictive models for patient outcomes, it reduces the complexity of processing high-dimensional data, such as medical tests and demographics, making the model more efficient and computationally cost-effective for large datasets and real-time predictions. We analyzed the time complexity of each attention mechanism to shed some light on how healthcare research can integrate attention-driven component based on available resources in Table 5.
@@
Table 5. Time complexities of various attention mechanisms in cNLP, from least to most efficient. Here, $N$ is the sequence length, $d$ is the embedding dimension, $M$ is the cross-attended length, $w$ is a local window size, $g$ is the number of global tokens, and $r$ is a low-rank dimension. “KG overhead” and $\epsilon$ indicate extra constant factors due to knowledge bases or entity-related modules. These values are based on official reports or estimated according to functionality.